# Mixture-of-RAG: Integrating Text and Tables with Large Language Models

### v2

@@ 4.3 Ablation Study
Figure 4: Scalability Analysis of Retrieval Performance across Different Corpus Sizes.
@@
We evaluate the scalability of MixRAG by comparing its retrieval performance across different document corpus sizes. As shown in Figure 4, MixRAG consistently maintains the highest HiT@K scores (K=1, 3, 5, 10) across all corpus scales, demonstrating its ability to preserve retrieval precision even as the corpus size increases. While baseline methods experience significant drops in performance as the document corpus increases, MixRAG exhibits the slowest rate of degradation, particularly in HiT@10, where its performance remains notably stable even with a larger document corpus. The retrieval stability reflects the scalability of MixRAG in handling growing document complexity. These results confirm that MixRAG effectively maintains retrieval precision even with larger and more complex document sets, showcasing its superior scalability compared to existing methods.
5 Conclusion and Limitations
@@
Conclusion. In this paper, we propose MixRAG, a framework designed for Heterogeneous Document RAG. To support this task, we introduce DocRAGLib, a benchmark dataset containing over 2000 heterogeneous documents and over 4000 QA pairs. MixRAG integrates hierarchical table summarization, cross-modal retrieval with LLM-based reranking, and RECAP-guided multi-step reasoning to support RAG over heterogeneous documents. Extensive experiments show that MixRAG outperforms existing baselines in retrieval precision and QA accuracy, establishing new state-of-the-art results for mixed-modality document grounding.
Limitations. Our Heterogeneous Document RAG task setup selects the top-1 retrieved document, which may limit performance on tasks requiring evidence aggregation from multiple document sources. In the future, extending the framework to incorporate multi-document retrieval and integration could better capture distributed evidence and support complex scenarios. In addition, MixRAG specializes in heterogeneous documents integration. It does not handle visual modalities (images, figures, or schematics) commonly found in real-world documents. Incorporating visual modalities into the RAG framework presents a promising direction for future research.
@@ Appendix A Dataset
A.1 Data Collection
The documents of DocRAGLib are derived from two sources: the existing public single-document QA dataset and web-scraped text-table heterogeneous data. First, we collect information from the publicly available single-document QA dataset MultiHiertt [30], which is designed for single-document QA tasks. Each document in MultiHiertt comprises textual content and hierarchical tables, with an average of 3.89 tables per document. In addition, we scrape contextual information related to the tables by utilizing the webpage links provided in another available single-table QA dataset HiTab [9], and then reconstruct heterogeneous documents containing both text and tables.
A.2 Data Cleaning
@@
Figure 5: Pipeline for ambiguity resolution during QA pair construction
@@
A.3 Data Statistics
@@
Table 4: Statistical Analysis of the DocRAGLib Dataset
@@
    Documents with >2 Tables 1,932 (88.71%)
@@
    Numerical Values >10,000 865 (19.36%)
    Numerical Values >100,000 181 (4.05%)
Appendix B Different Table Representation Strategies
To understand how summarization granularity affects information quality and retrieval, we compare three table summarization strategies: Table-Level, Row Column Level (RCL), and Cell Level.
Figure 6: An Example of Three Level Table Summarization
Figure 6 shows three summarization strategies applied to the same table. As illustrated, table-level summaries offer an overview but lose fine-grained structural information. The over-abstraction reduces the discriminative power during retrieval, leading to decreased precision and increased ambiguity when distinguishing between similar tables. In addition, cell-level summaries provide highly specific and detailed descriptions, capturing cell-level semantics that are often useful for retrieval. However, this fine granularity comes at the cost of redundancy and low retrieval efficiency, as repeated expressions across cells inflate input length and dilute relevance during retrieval. We do not include cell-level summarization in our experimental design due to its high cost and significant computational overhead.
Therefore, we focus on row and column level (RCL) summarization, which strikes a balance between structure preservation and expression compactness by summarizing content at the header level. To further reduce noise, we additionally mask out non-semantic numeric values in table cells. The RCL middle-ground representation retains the hierarchical structure of tables and preserves header-level semantics essential for retrieval. This design enables more effective alignment between table content and question, serving as a foundation for retrieval in heterogeneous documents.
Appendix C Baselines
This section provides detailed descriptions of the baseline methods used for both retrieval and QA tasks.
For the retrieval task, we include several representative methods:
  • LangChain: A semi-structured retrieval pipeline that processes tabular and textual data by summarizing content with LLMs and conducting similarity-based retrieval.
  • SelfRAG [1]: A self-reflective retrieval framework that dynamically retrieves passages and uses reflection tokens to iteratively assess and refine outputs based on the model’s own feedback.
  • Table Retrieval [4]: A method tailored for table retrieval that discovers join relationships and applies mixed-integer programming-based re-ranking to optimize both table-query and table-table relevance.
For the QA task, we consider several widely used reasoning frameworks that have demonstrated strong performance in answering structured and semi-structured questions:
  • COT [25]: A reasoning-based prompting strategy that guides the model to generate intermediate reasoning steps, improving its ability to solve complex problems.
  • EEDP [19]: A multi-step reasoning framework that decomposes the answer process into explanation, refinement, and decision phases to enhance accuracy.
  • POT [6]: A technique that prompts the LLM to produce a structured program for reasoning, which is then executed externally to yield precise results, especially for numerical problems.
Appendix D Evaluation Metrics
@@
    $\text{HiT@K}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}\left[d_{i}^{\ast}\in\text{TopK}(q_{i})\right],$ (11)
@@
    $\text{EM}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}\left[a_{i}=\hat{a}_{i}\right],$ (12)
@@
Appendix E Algorithm Pseudo Code of MixRAG
Algorithm 1 illustrates the overall design of MixRAG (excluding heterogeneous document representation) in three sequential stages. (1) Ensemble Retrieval: BM25 and vector-based retrieval run in parallel on the query $Q$, producing $n$ candidates from the table-level corpus and $m$ candidates from the row–column-level corpus, which are then merged into an initial set of $K$ candidate documents. (2) LLM-based Retrieval: For the $K$ candidate documents, we first retain the passages most semantically similar to the question and reconstruct them in their original order into “new documents,” then feed all new documents into the LLM to accurately select the target document $D^{*}$. (3) QA Inference via RECAP: In the RECAP stage, the LLM restates the question, extracts evidence, and produces a computation formula for an external calculator to execute; a rule-based mechanism then selects the final answer $A$ from the LLM’s output and the calculator’s result.
Algorithm 1 MixRAG: Retrieval-Augmented Generation Workflow
Data: $Q$: question, $C_{\mathrm{BM25}}$: BM25 corpus, $C_{\mathrm{Embed}}$: embedding corpus, BM25: BM25 retriever, Embed: dense embedding model, LLM: large language model, Calc: external calculator for executing formulas provided by RECAP, $n$: number of BM25 candidates; $m$: number of embedding candidates; $K$: final candidate set size; $\Theta$: chunk-filter threshold
Result: $A$: predicted answer to the question
1:function MixRAG($Q,\,C_{\mathrm{BM25}},\,C_{\mathrm{H\text{-}RCL}},\,\text{BM25},\,\text{Embed},\,\text{LLM},\,\text{Calc},\,n,\,m,\,K,\,\Theta$)
2:  // Stage 1: Ensemble Retrieval
3:  $S_{1}\leftarrow$ BM25Retrieve($\text{BM25},\,C_{\mathrm{BM25}},\,Q,\,n$)
4:  $S_{2}\leftarrow$ EmbedRetrieve($\text{Embed},\,C_{\mathrm{Embed}},\,Q,\,m$)
5:  $S\leftarrow$ MergeCandidates($S_{1},\,S_{2},\,K$)
6:  // Stage 2: LLM-based Retrieval
7:  $D^{*}\leftarrow$ LLMBasedRetrieve($\text{LLM},\,Q,\,S,\,\Theta$)
8:  // Stage 3: QA Inference (RECAP)
9:  $A\leftarrow$ RECAP($\text{LLM},\,\text{Calc},\,Q,\,D^{*}$)
10:  return AA
11:end function
12:function BM25Retrieve($\text{BM25},\,\text{corpus},\,Q,\,n$)
13:  return top-$n$ documents by BM25 score
14:end function
15:function EmbedRetrieve($\text{Embed},\,\text{corpus},\,Q,\,m$)
16:  return top-$m$ documents by $\cos\bigl{(}\text{Embed}(Q),\,\text{Embed}(\text{chunk})\bigr{)}$
17:end function
18:function MergeCandidates($S_{1},\,S_{2},\,K$)
19:  $U\leftarrow S_{1}\cup S_{2}$ $\triangleright$ remove duplicates
20:  return top-$K$ of $U$
21:end function
22:function LLMBasedRetrieve($\text{LLM},\,Q,\,S,\,\Theta$)
23:  $S^{\prime}\leftarrow\emptyset$
24:  for all $D\in S$ do
25:   $M\leftarrow$ SelectTopChunks($Q,\,D,\,\Theta$) $\triangleright$ retain chunks most relevant to the question
26:   $D^{\prime}\leftarrow$ Recombine(M,DM,\,D)
27:   $S^{\prime}\leftarrow S^{\prime}\cup\{D^{\prime}\}$
28:  end for
[…]
30:end function
31:function RECAP($\text{LLM},\,\text{Calc},\,Q,\,D^{*}$)
32:  $r\leftarrow\text{LLM.Restate}(Q)$ $\triangleright$ (1) Analyze Question Type
33:  $e\leftarrow\text{LLM.Extract}(r,\,D^{*})$ $\triangleright$ (2) Extract Relevant Data
34:  $\ell\leftarrow\text{LLM.Compute}(e)$ $\triangleright$ (3) Perform Logical Reasoning
35:  $a_{\mathrm{LLM}}\leftarrow\text{LLM.Answer}(\ell)$ $\triangleright$ (4) Generate LLM Answer
36:  $f\leftarrow\text{LLM.PresentFormula}(\ell)$
37:  $a_{\mathrm{Calc}}\leftarrow\text{Calc.Evaluate}(f)$ $\triangleright$ (5) Generate Calc Answer
38:  return SelectFinal($a_{\mathrm{LLM}},\,a_{\mathrm{Calc}},\,Q$) $\triangleright$ (6) Provide Final Answer
39:end function
Appendix F Hyperparameter Experiment
@@
Table 5: Retrieval performance under different $(n,m)$ configurations with deduplicated document count $k$.
@@
Appendix G Comparison of RECAP with Various Baselines
The experimental results of RECAP and baseline methods on the DocRAGLib dataset are shown in Table 6. Five LLMs are used in the experiments, with the EM metric as the evaluation criterion. To ensure fairness in evaluation, we strictly constrain the output formats of the LLMs and tailor the EM-based evaluation mechanism to the different answer types in the dataset.
Among the baseline methods, PoT performs the best, mainly due to its program execution logic, which improves the accuracy of calculation-based questions in the DocRAGLib dataset. However, RECAP outperforms PoT in EM across all five LLMs, demonstrating its superiority. RECAP improves the accuracy of calculation-based questions through reasoning via external calculation functions, reducing output instability compared to program-generated methods like PoT. Furthermore, GPT-4o shows superior performance with the RECAP method (EM = 0.6466) compared to the other four models and is ultimately selected for the Multi-Step Reasoning Module in MixRAG.
Table 6: Comparison of RECAP in MixRAG Document-QA Performance with Various Baselines
Models
Method GPT-4o GPT-4o mini Qwen2.5-32B Mistral-Nemo-2407 LLama3.1-7B
Direct 0.3534 0.2355 0.2180 0.1003 0.1579
CoT 0.4687 0.3684 0.4837 0.2506 0.2531
EEDP 0.4737 0.3759 0.4862 0.2757 0.2080
PoT 0.6190 0.5564 0.5589 0.3409 0.2932
RECAP 0.6466 0.5815 0.6416 0.4411 0.4060
Appendix H Evaluation of the Reranking Performance of Different Models
In contrast, the cross-encoder models, including ms-marco-MiniLM-L12-v2, ms-marco-MiniLM-L6-v2, and ms-marco-MiniLM-L2-v2, perform fine-grained token-level interactions via full-attention between query and document, yielding highly calibrated relevance scores at the expense of greater computational cost. Finally, we include the BGE reranker series (e.g., bge-reranker-v2-m3, bge-base-en-v1.5, bge-large-en-v1.5, and bge-base-en) to represent bi-encoder architectures that first encode query and document separately into fixed embeddings and then score relevance via cosine similarity. Together, these three classes span the methodological spectrum from generative inference to cross-attention scoring and pure embedding matching.
Table 7: Evaluation of the Reranking Performance of Different Models
Rerank Method Model HiT@1 HiT@3 HiT@5

### v3

@@ 5.4. Scalability Analysis
Figure 5. Scalability Analysis of Retrieval Performance across Different Corpus Sizes.
@@
We evaluate the scalability of MixRAG by comparing its retrieval performance with several baselines across different document corpus sizes in GPT-4o. Standard RAG is excluded from this analysis due to its consistently low retrieval accuracy. As shown in Figure 5, MixRAG consistently maintains the highest HiT@K scores (K=1, 3, 5, 10) across all corpus scales, demonstrating its ability to preserve retrieval precision even as the corpus size increases. While baseline methods experience significant drops in performance as the document corpus increases, MixRAG exhibits the slowest rate of degradation, particularly in HiT@10, where its performance remains notably stable even with a larger document corpus. The retrieval stability reflects the scalability of MixRAG in handling growing document complexity. These results confirm that MixRAG effectively maintains retrieval precision even with larger and more complex document sets, showcasing its superior scalability compared to existing methods.
5.5. Failure Case Study
In this section, we analyze the failure cases in our experiments. We categorize these failure cases into three main types.
Retrieval Failure Due to Incomplete Question. When questions are unclear or lack explicit references to key information within the documents, MixRAG struggles to identify content that is strongly related to the question, leading to retrieval wrong document. This problem is particularly common in RAG systems, as they heavily rely on explicit question-document relationships. Without clear references in the question to guide the retrieval, it may retrieval irrelevant document, which in turn impacts the accuracy of the subsequent task.
Reasoning Failure Due to Complex Multi-Step Tasks. When the question requires complex multi-step reasoning beyond the LLM’s capabilities in a single prompt, the LLM may fail to generate the correct answer. It is common in tasks involving numerical calculations and logical deductions.
Reasoning Failure Due to Ambiguity in Reasoning Paths. When questions have multiple valid interpretations, our Multi-Step Reasoning Module may choose an incorrect reasoning path, resulting in errors in the final answer. This issue is common in LLM-based methods, particularly when tasks have multiple possible answers or reasoning approaches.
In our future work, we plan to design a question decomposition approach to break down the complex questions into sub-questions, improving retrieval accuracy by ensuring the clarity of each sub-question. Additionally, the decomposition will explicitly display the reasoning paths for LLM execution, supporting multi-step reasoning through multiple calls, which enhances the LLM’s ability to handle complex tasks. Furthermore, we aim to introduce validation for reasoning paths to reduce ambiguity and guarantee the selection of the most appropriate path, ultimately improving the accuracy and reliability of the reasoning process. By addressing these failure cases and incorporating these improvements, we aim to improve the performance and robustness of MixRAG in future work.
6. Conclusion
@@
In this paper, we investigate the challenge of question answering over heterogeneous documents. To support this task, we introduce DocRAGLib, a benchmark dataset containing over 2,000 heterogeneous documents and over 4,000 QA pairs. To effectively address the challenges in this task, we propose the MixRAG framework. MixRAG integrates hierarchical table summarization, cross-modal retrieval with LLM-based reranking, and RECAP-guided multi-step reasoning to support RAG over heterogeneous documents. Extensive experiments show that MixRAG outperforms existing baselines in retrieval precision and QA accuracy, establishing new state-of-the-art results for heterogeneous document grounding while maintaining a superior trade-off between effectiveness and efficiency.
Acknowledgements.
We would like to thank Haoyu Dong for his valuable assistance and constructive suggestions, which greatly improved this work.
@@ Appendix A Dataset
A.1. Data Cleaning
@@
Figure 6. Pipeline for ambiguity resolution during QA pair construction
@@
In addition to the automated cleaning pipeline described above, we conducted a manual evaluation stage to further enhance data quality. Evaluation team members reviewed a random sample of the dataset; each member examined at most 100 heterogeneous documents together with all of their associated QA pairs. This manual audit led to the removal of 249 documents and 370 QA pairs. The principal reasons for filtering were: (1) ambiguous question formulations (in these cases only the question itself was discarded); (2) mismatches between textual passages and table information; (3) missing or incomplete textual content; and (4) irregular table structures (e.g., merged cells that disrupt the hierarchical layout). For the latter three issues, because the documents contained irreparable defects, we deleted the entire document along with all related questions. This additional manual quality-control step markedly improved the dataset’s overall accuracy, consistency, and readability.
A.2. Data Statistics
@@
Table 6. Statistical Analysis of the DocRAGLib Dataset
@@
    Documents with ¿2 Tables 1,932 (88.71%)
@@
    Numerical Values ¿10,000 865 (19.36%)
    Numerical Values ¿100,000 181 (4.05%)
Appendix B Evaluation Metrics
@@
    (11) $\text{HiT@K}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}\left[d_{i}^{\ast}\in\text{TopK}(q_{i})\right],$
@@
    (12) $\text{EM}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}\left[a_{i}=\hat{a}_{i}\right],$
@@
  • Token Usage: We measure the computational cost via the number of tokens consumed by the LLM. This metric is decomposed into preprocessing tokens and generation tokens.
  • Time: We record the average end-to-end wall-clock time (in seconds) required to process a query.
Appendix C Hyperparameter Experiment
@@
Table 7. Retrieval performance under different $(n,m)$ configurations with deduplicated document count $k$.
@@
Table 8. Evaluation of the Reranking Performance of Different Models
Reranker Model HiT@1 HiT@3 HiT@5